OpenStudy (anonymous):

Alright, Scenario: Rhianna says she can draw different functions that have the same x-intercepts and the same domain and range. Her teammates say "No, that's impossible!" But Rhianna insists,"It is possible if we just try to sketch the graphs." a. What if the x-intercepts are (-5,0),(2,0), and (6,0), the domain is −5≤x≤ and the range is −4≤y≤10 Is there more than one function possible? Give examples.

OpenStudy (anonymous):

I don't understand any of it at all.

OpenStudy (baldymcgee6):

do you havr the link?

OpenStudy (baldymcgee6):

there are 3 diff x intercept when the question says they are the same.. doesnt make sense

OpenStudy (anonymous):

It isn't in the program. They skip it Here's the link to where it should be: http://www.cpm.org/students/homework/A2C_Problems/A2C_Ch1_Answers/A2C_hwk_ch1.html It's 1-46 But its not on there.

OpenStudy (baldymcgee6):

maybe because the question doesnt make sense.. lol

OpenStudy (anonymous):

This is what it says word for word in the book: a. What if the x-intercepts are (-5,0) , (2,0) , and (6,0), The domain is:\[-5 \le x \le 7\]and the range is:\[-4 \le y \le 10\] Is more than one function possible? Give examples to help explain why or why not.

OpenStudy (baldymcgee6):

more than one function is totally possible, but they dont have the same x intercepts like it said earlier.

OpenStudy (anonymous):

So the answer would be no then, because the x intercepts aren't the same? Because in the scenario it states they have to be that same. Thats fine, moving on. Part b. b. What if the x-intercepts are (-4,0) and (2,0) the domain is all real numbers, and the range is \[y \ge -8\] Is there more then one function possible? Give examples of multiple functions or explain why there can only be one.

OpenStudy (baldymcgee6):

there will always be more than one function possible

OpenStudy (anonymous):

What are some exmaples

OpenStudy (baldymcgee6):

draw or function?

OpenStudy (anonymous):

Functions,

OpenStudy (baldymcgee6):

y = x^2 -8

OpenStudy (anonymous):

One more

OpenStudy (baldymcgee6):

man you're needy ;) y = (1/9)x^2 -8

OpenStudy (baldymcgee6):

please throw another medal my way.. :)